OpenStudy (anonymous):

3b-4 -------=c , for b 2

OpenStudy (anonymous):

\[\frac{ 3b-4 }{ 2 }=c\] You can multiply everything by 2 to get rid of the fraction.

OpenStudy (anonymous):

You will get 3b-4=2c

OpenStudy (anonymous):

From here just solve for b

OpenStudy (anonymous):

how does multiplying by two get rid of the fraction????

OpenStudy (anonymous):

It like you were multiplying 2/1*1/2 You'll get 2/2 which is 1. In other words, they cancel out.

OpenStudy (anonymous):



OpenStudy (anonymous):

ok thank you very much! :)

OpenStudy (anonymous):

After solving for b. You should get \[b=\frac{ 2c+4 }{ 3 }\]